OpenStudy (anonymous):

Solve for x if e^2x+1=5

OpenStudy (luigi0210):

Get e by itself first, so move that 1

OpenStudy (luigi0210):

Bro, are you dead or something?

OpenStudy (anonymous):

Sorry, I trying to figure this out.

OpenStudy (anonymous):

lol and I am not a guy ln(e^2x+1)=ln(5)

OpenStudy (psymon):

do what luigi suggested first.

OpenStudy (anonymous):

cross out the ln and e

hartnn (hartnn):

is it e^(2x+1) or e^(2x) +1 ??

hartnn (hartnn):

is the 1 included in the exponent ?

OpenStudy (anonymous):

Yes it is the 2x+1 is the exponent together

OpenStudy (psymon):

Oh x_x My apologies then.

hartnn (hartnn):

cool, so take ln on both sides,

hartnn (hartnn):

\(e^a=b \implies a = \ln b\)

OpenStudy (anonymous):

I am not sure what b and a is?

hartnn (hartnn):

nevermind, take ln on both sides of e^(2x+1)=5 what do u get ?

OpenStudy (anonymous):

2x +1= ln5

hartnn (hartnn):

just subtract 1 from both sides

hartnn (hartnn):

then divide both sides by 2, and you would get 'x'

OpenStudy (lasttccasey):

Not to steal the answer but sometimes seeing it like this helps, good job explaining @hartnn e^(2x+1)=5 Original Question ln(e^(2x+1)=ln(5) Take the natural log of both sides (ln) 2x+1=ln(5) The ln cancels the e and the right side doesn't change 2x=ln(5)-1 subtract the one over and then divide by two to get:\[x=\frac{ \ln (5)-1 }{ 2 }\]

OpenStudy (anonymous):

Thanks @hartnn and @lasttccasey very clear. But I have one question about the part where ln and e getting canceled? How does that get canceled.

OpenStudy (lasttccasey):

ln is the opposite of e functions.

hartnn (hartnn):

\(\Large \ln e^y = y \ln e = y\times 1\) as ln e =1

hartnn (hartnn):

by the property that \(\log_aa=1\)

OpenStudy (anonymous):

oh so it = 1 that is why it is canceled.

OpenStudy (lasttccasey):

ln e does = 1 because its actually ln(e^1) so the power on the e becomes the remainder. log and natural logs are slightly different.

OpenStudy (anonymous):

I see, wow thanks, the rest is clear.

OpenStudy (lasttccasey):

You're welcome.